jenny and penny brought a llama for £4500. jenny paid £1215 towards this total. what % of the total cost did jenny pay
Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

jenny and penny brought a llama for £4500. jenny paid £1215 towards this total. what % of the total cost did jenny pay
----
Ans: x/100 = 1215/4500
x = 121500/4500
---
x = 27
----
Ans: 27%
------------
Cheers,
Stan H.
-----------

Answer by addingup(3677)

(1215/4500)*100= percentage Jenny paid. You punch the numbers in your calculator, first divide and then multiply the result times 100.